Betty has 10 more dimes than quarters. If she has $3.45, how many coins does she have?
Answer by mananth(16949)

let her haave x quarters
she will have x+10 dimes
25x+10(x+10)= 345 cents
25x+10x+100=345
35x=345-100
35x=245
x= 7 ------ quarters
y= x+10 = 7+10 = 17 dimes